OpenStudy (anonymous):

What is the equation of this line in standard form using integer coefficients?

OpenStudy (anonymous):

Are you going to post the equation?

OpenStudy (anonymous):

\[y=\frac{ 5 }{ 8 }x-9\]

OpenStudy (ash2326):

@alexawb multiply both sides by 8, could you do that?

OpenStudy (ash2326):

\[y=\frac 5 8 x-9\] \[8\times y=8\times (\frac 5 8 x-9)\] could you take it from here?

OpenStudy (anonymous):

Yess thank you!

OpenStudy (ash2326):

After you're done, just post the simplified equation here.

OpenStudy (anonymous):

y=5x-72 ? But is that in the integer coefficient form?

OpenStudy (ash2326):

Yes, but we need to move things a little, subtract both sides by y and add 72 to both the sides. Would you do that?

OpenStudy (anonymous):

I got it, 5x + 8y = -72 :) Thanks for your help

OpenStudy (ash2326):

You're welcome:)